OpenStudy (vera_ewing):

It takes 250.0 J to raise the temperature of a 50.0 g sample of a metal by 10.0C. What is the specific heat capacity of this metal? A. 1.25 x 105 J/g K B. 5.00 x 10-4 J/g K C. 50.0 J/g K D. 1.25 x 102 kJ/g K E. 0.500 J/g K

OpenStudy (vera_ewing):

@sweetburger please help!

OpenStudy (sweetburger):

q=MC Delta T

OpenStudy (sweetburger):

so 250J =q 50g = m and 10C is equal to Delta T

OpenStudy (sweetburger):

250J = 50g(C)(10C) = 250/(500) = 1/2 or .5

OpenStudy (vera_ewing):

So what would the answer be?

OpenStudy (vera_ewing):

E?

OpenStudy (sweetburger):

yes